OpenStudy (anonymous):

Find the greatest common factor of 15x 2 y 3 and -20x 3 yz.

OpenStudy (anonymous):

there's no gcf

OpenStudy (jdoe0001):

hint: \(\begin{array}{cccllll} 15x^2y^3&&-20x^3yz\\ \uparrow &\qquad &\uparrow \\ 3\cdot {\color{brown}{ 5}}\cdot {\color{brown}{ x\cdot x}}\cdot y\cdot {\color{brown}{ y}}\cdot y&&-2\cdot 2\cdot {\color{brown}{ 5}}\cdot x\cdot {\color{brown}{ x\cdot x\cdot y}}z \end{array}\)

OpenStudy (anonymous):

@jdoe0001 there's not gcf bc it also has 2 n 2 can't go into 15 evenly n 2 is a prime # so nun else can be divided into it bc of tht